The length of a new rectangular playing field is 6 yards longer than triple the width. If the perimeter of the rectangular playing field is 492 yards. What are the dimensions?
Answer by rfer(12730)

2x+2(3x+6)=492
2x+6x+12=492
8x=480
x=60 yds
3x+6=186 yds